# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 1.5.4 Applied category theory
The compositional nature of EPHS can be made precise using category theory. Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics Mac Lane (1998). In recent decades, category theory has also grown beyond pure math to organize scientific and engineering disciplines as well, giving birth to ‘applied category theory’ Fong and Spivak (2019). A specific approach to categorical systems theory Myers (2022) that has inspired this work is based on the following principles, which can essentially also be found in the work of Willems Willems (2007).
@@
    A shift of focus from individual systems to how systems relate to each other.
@@
    The description of systems involves syntax and semantics. Syntax typically consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically. Additionally, the syntax can be displayed and manipulated diagrammatically as well as in text. An explicit understanding of the translation from syntax to semantics serves as a guide for computer implementation.
For instance, there is a line of research which formalizes graphical languages for systems and process such as block diagrams Baez and Erberle (2015), Petri nets Baez and Pollard (2017), junctions of bond graphs Coya (2017), and passive linear circuits Baez and Fong (2018). These works consider a graphical syntax as some sort of string diagrams, established via a generators and relations approach.
More closely related to our work, another line of research uses the theory of operads to define a graphical syntax which supports hierarchical (de-)composition, see Baez et al. (2020); Foley et al. (2021). Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system), see Leinster (2004); Yau (2016). As proposed in Lohmayer and Leyendecker (2022), the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in Spivak (2013), together with its relational semantics used to express database queries, see also Spivak (2014); Yau (2018); Breiner et al. (2020). In a similar lane, directed wiring diagrams are used to formalize nested systems of Moore machines and ODEs with outputs, see Vagner et al. (2015); Schultz et al. (2019). A combination of dynamical systems based on directed and undirected wiring diagrams is presented in Libkind (2020); Libkind et al. (2021). The composition of lossless port-Hamiltonian systems was recently studied from a categorical perspective in Lynch (2022).
@@ 1.6 Outline
Section 2 defines system interfaces. Based on this, Section 3 defines interconnection patterns. Section 4 shows how every interconnection pattern implies a relation between the port variables of its interfaces. Section 5 discusses how exergy quantifies the energy stored in a system that is available for doing work. Based on this, Section 6 introduces the primitive systems. Section 7 defines composite systems and how they are formed using the graphical syntax. Section 8 shows that EPHS models are thermodynamically consistent. Section 9 presents an example of a motor model, together with a comparison to a classical bond graph. Section 10 concludes with a discussion.
@@ 2.1 Namespaces
Given two names $p$ and $n$, we say that $p$ is a prefix of $n$, and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$. In case $s\neq[]$, we say that $p$ is a strict prefix of $n$ and write $p<n$. ∎
@@
It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
@@
A set of names $N$ is prefix-free if for all $n,n^{\prime}\in N$, $n$ is not a strict prefix of $n^{\prime}$. A namespace is a finite, prefix-free set of names. Let $\mathbf{Namespace}$ refer to the set of all namespaces. ∎
@@
Sets of names are equivalent to prefix trees, called tries in computer science Fredkin (1960).
@@
Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names, which is not prefix-free, at least one name corresponds to an internal node. In the following, we think of namespaces both as sets of names and as tries.
@@
Given a namespace $N$, and a function $f\colon N\to\mathbf{Namespace}$, the named sum of $f$ is
@@
Given $N=\{\mathtt{q},\,\mathtt{p.1},\,\mathtt{p.2}\}$, we have
$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{s.t}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.q},\,\mathtt{l.p.1},\,\mathtt{l.p.2},\,\mathtt{s.t.q},\,\mathtt{s.t.p.1},\,\mathtt{s.t.p.2}\bigr{\}}\,.\qed$
@@
Given a namespace $N$ and a function $f\colon N\to\mathbf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say it is prefix-free. ∎
@@
The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted on top. ∎
Namespaces and named sums provide a simple and effective formalism for working with large hierarchically-structured systems.
2.2 Packages
An interface has a namespace, where we interpret each name as referring to a port. However, this is not the only ingredient to an interface, as each port has two associated attributes. This motivates the definition of packages:
@@ 2.2 Packages
Given a set $A$, we define
$\mathbf{Pkg}(A)\>=\>\Bigl{\{}(N,\,\tau)\,\mid\,N\in\mathbf{Namespace},\,\tau\colon N\to A\Bigr{\}}\,.$
An element $(N,\,\tau)\in\mathbf{Pkg}(A)$ is called a package of elements of $A$. We also just write $\tau\in\mathbf{Pkg}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
The name ‘package’ is drawn from software engineering; one might consider a software package to be a ‘package of types and procedures’.
A package $(N,\,\tau)\in\mathbf{Pkg}(A)$ can be seen as a trie $N$ with an element $\tau(n)\in A$ associated to each leaf node $n\in N$. Hence, the concept of named sum carries over from $\mathbf{Namespace}$ to $\mathbf{Pkg}(A)$, for a fixed set $A$. Named sum can then be seen as a flattening operation, analogous to flattening a list of lists. Specifically, for the set of all lists of elements of $A$
$\operatorname{\mathbf{List}}(A)\>=\>\bigl{\{}(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\bigr{\}}\,,$
flattening is a map $\operatorname{\mathbf{List}}(\operatorname{\mathbf{List}}(A))\to\operatorname{\mathbf{List}}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ]. Similarly, named sum is a map $\mathbf{Pkg}(\mathbf{Pkg}(A))$ to $\mathbf{Pkg}(A)$. An input to this is a trie $N$, which has associated to each leaf node $n\in N$ another trie $N_{n}$, whose leaves carry elements of $A$. The named sum operation then simply grafts onto each node $n\in N$ its associated trie $N_{n}$, carrying along the elements of $A$ at the leaves. In terms of category theory, the flattening operation makes $\operatorname{\mathbf{List}}$ and $\mathbf{Pkg}$ examples of monads.
@@ 2.3 Bundle of quantities
The first attribute associated to any port represents the physical quantity that can be exchanged via the port, e.g. entropy or momentum, along with a state space in which values of the quantity live. For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidian space are considered, etc.
@@
To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see Pavelka et al. (2014); Lohmayer et al. (2021) for details.
@@
Let the parity of each quantity with respect to time-reversal transformation be given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
@@
A comprehensive collection of physical quantities with their associated parities would be a built-in feature of a software implementation.
2.4 Interfaces as packages
Each port has two attributes, namely the already discussed quantity and further a Boolean attribute that indicates whether the port is a power port or a state port. Only power ports allow for exchange of energy associated with the attributed quantity. A state port is required whenever the behavior of a system depends on the state of another system, although no energy is exchanged between the two.
@@ 2.4 Interfaces as packages
An interface is a package of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
$\mathbf{Interface}\>=\>\mathbf{Pkg}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.\qed$
@@
Let $I=(N,\,\tau)$ be an interface with namespace $N=\{\mathtt{p_{1}},\,\mathtt{p_{2}},\,\mathtt{q}\}$ and attributes given by $\tau(\mathtt{p_{1}})=\tau(\mathtt{p_{2}})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$ and $\tau(\mathtt{q})=\left((\mathbb{R},\,\mathtt{displacement}),\,\mathsf{s}\right)$. Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy, along with information about the current momentum. Port $\mathtt{q}$ is a state port that can only exchange information about the current displacement. ∎
We note that the named sum of interfaces is a function $\mathbf{Pkg}(\mathbf{Interface})\to\mathbf{Interface}$.
@@ 2.5 Bundles of port variables
A state port has one port variable called state. A power port has two further port variables called flow and effort. Given an interface, the port variables of all its ports combined together take values in a vector bundle, called the bundle of port variables associated to the interface.
@@
The state space associated to interface $I$ is the Cartesian product
$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.\qed$
@@
Given any interface $I$, each port $p\in I$ has an associated state variable $p\mathtt{.x}\in\mathcal{X}_{p}$. If $p$ is a power port, it additionally has a flow variable $p\mathtt{.f}$ and an effort variable $p\mathtt{.e}$. All port variables implicitly depend on time. As detailed in Section 6.1, the flow variable is the instantaneous rate of change of the state variable and the effort variable is the differential of a function of the state variable. Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see e.g. Lee (2012)). The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\in\mathbb{R}$ yields the instantaneous power that is exchanged through port $p$.
Now we can define the bundle of port variables associated to an interface:
@@
Let $I$ be an interface. Based on the Boolean attribute of its ports, we can write its associated state space as the Cartesian product $\mathcal{X}_{I}\cong\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports. Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote all flow and effort variables. The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$. Hence, $\mathcal{P}_{I}\cong\mathbb{T}\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$.
{$\begin{tikzcd}\qed$
@@
Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has total space (or port space)
@@
and base space (or state space)
@@
3 Interconnection patterns
As introduced in Section 1.2, an interconnection pattern expresses how a composite system can be formed from a finite number of subsystems. The syntax defines the entirety of such expressions and their composition, i.e. their hierarchical nesting.

### v2

@@ 1.5.4 Applied category theory
Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics (see, e.g., [10]). In recent decades, category theory has grown beyond pure math to organize scientific and engineering disciplines, giving birth to applied category theory (see, e.g. [9]). The following ideas from categorical systems theory (see, e.g., [19]) have inspired our work.
@@
    The shift of focus from individual systems to how systems relate to each other.
@@
    The description of systems involves syntax and semantics. Syntax consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically. An explicit understanding of the translation from syntax to semantics serves as a guide for computer implementation.
A prominent line of research formalizes graphical languages such as block diagrams [17], Petri nets [31], junctions of bond graphs [32], and passive linear circuits [33]. In each case, the graphical syntax is defined as some sort of string diagrams, using a generators and relations approach.
Another line of research uses the theory of operads to define a graphical syntax that supports hierarchical (de-)composition (see, e.g., [34, 35]). Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system) (see, e.g., [36, 37]). As proposed in [38], the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in [16], together with a relational semantics used to express database queries, see also [39, 40]. Directed wiring diagrams have been used to formalize nested systems of Moore machines and ODEs with inputs and outputs [41, 42]. A framework for discrete and continuous dynamical systems with directed and undirected notions of composition is presented in [43]. The composition of lossless PHS is studied from a categorical perspective in [44].
@@ 1.6 Outline
Section 2 defines system interfaces. Based on this, Section 3 defines interconnection patterns. Section 4 shows how an interconnection pattern implies a relation between the port variables of its interfaces. Section 5 discusses how exergy quantifies the energy stored in a system that is available for doing work. Based on this, Section 6 introduces the primitive systems. Section 7 defines composite systems and how they are formed using the graphical syntax. Section 8 shows that EPHS models are thermodynamically consistent. Section 9 presents an example of a direct current (DC) motor model, along with a comparison to a classical bond graph. Section 10 concludes with a discussion.
@@ 2.1 Namespaces
Given two names $p$ and $n$, we say that $p$ is a prefix of $n$ and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$. In case $s\neq[]$, we say that $p$ is a strict prefix of $n$ and write $p<n$. ∎
@@
It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also, $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
@@
A set of names $N$ is prefix-free if for all $n,n^{\prime}\in N$, $n$ is not a strict prefix of $n^{\prime}$. A namespace is a finite, prefix-free set of names. Let $\mathsf{Namespace}$ refer to the set of all namespaces. ∎
@@
Sets of names are equivalent to prefix trees, called tries in computer science [45].
@@
Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names that is not prefix-free, at least one name corresponds to an internal node. In the following, we think of namespaces both as sets of names and as tries.
@@
Given a namespace $N$, and a function $f\colon N\to\mathsf{Namespace}$, the named sum of $f$ is
@@
Given $N=\{\mathtt{p},\,\mathtt{q}\}$, we have
$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{r}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.p},\,\mathtt{l.q},\,\mathtt{r.p},\,\mathtt{r.q}\bigr{\}}\,.\qed$
@@
Given a namespace $N$ and a function $f\colon N\to\mathsf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say, it is prefix-free. ∎
@@
The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted onto it. ∎
As shown next, we can use namespaces and named sums to develop a simple and effective formalism for working with hierarchically-structured data.
2.2 Directories
The interface of a system is given by a namespace, where each name refers to a port. However, this is not all, as each port has two associated attributes. This motivates the definition of directories:
@@ 2.2 Directories
Given a set $A$ (of possible attributes), we define
$\mathsf{Dtry}(A)\>=\>\Bigl{\{}\,(N,\,\tau)\,\mid\,N\in\mathsf{Namespace},\,\tau\colon N\to A\,\Bigr{\}}\,.$
An element $(N,\,\tau)\in\mathsf{Dtry}(A)$ is called a directory of elements of $A$. We also just write $\tau\in\mathsf{Dtry}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
The name ‘directory’ points to the similarity with hierarchical file systems, where complete paths (names) are associated to file contents (attributes).
A directory $(N,\,\tau)\in\mathsf{Dtry}(A)$ can be seen as a trie $N$, where each leaf node $n\in N$ has an associated attribute $\tau(n)\in A$. Hence, the concept of named sum carries over from $\mathsf{Namespace}$ to $\mathsf{Dtry}(A)$, for any fixed set $A$. Named sum can then be seen as a flattening operation, analogous to flattening a list of lists. Specifically, for the set of all lists of elements of $A$
$\mathsf{List}(A)\>=\>\Bigl{\{}\,(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\,\Bigr{\}}\,,$
flattening is a map $\mathsf{List}(\mathsf{List}(A))\to\mathsf{List}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ]. Similarly, named sum is a map $\mathsf{Dtry}(\mathsf{Dtry}(A))$ to $\mathsf{Dtry}(A)$. An input to this is a trie $N$, which has associated to each leaf node $n\in N$ another trie $N_{n}$, whose leaves carry elements of $A$. The named sum operation then simply grafts onto each node $n\in N$ its associated trie $N_{n}$, carrying along the elements of $A$ at the leaves. In terms of category theory, the flattening operation, together with a unit operation, makes $\mathsf{List}$ and $\mathsf{Dtry}$ examples of monads.
@@ 2.3 Bundle of quantities
The first attribute associated to a port represents the physical quantity that can be exchanged via the port, e.g. momentum or entropy, along with a state space in which values of the quantity live. For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered, or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidean space are considered, etc.
@@
We note that $(\mathbb{R},\,\mathtt{energy})\notin\mathbf{Q}$, because in the context of EPHS, internal energy is a function of the state variable entropy, and not vice versa, see also [3].
To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see [46, 3] for details.
@@
The parity of each quantity with respect to time-reversal transformation is given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
@@
A comprehensive collection of physical quantities with their associated parities would be part of a software implementation.
2.4 Interfaces as directories
Each port has two attributes: the previously discussed quantity, and a Boolean attribute that indicates whether the port is a power port or a state port. Indeed, only power ports allow for an exchange of energy, based on an exchange of the respective quantity. State ports are required whenever the behavior of one system depends on the state of another system, although no energy is exchanged between the two.
@@ 2.4 Interfaces as directories
An interface is a directory of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
$\mathsf{Interface}\>=\>\mathsf{Dtry}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.$
Here, element $\mathsf{p}$ stands for power port and $\mathsf{s}$ stands for state port. ∎
@@
Let $I=(N,\,\tau)$ be an interface with namespace $N=\{\mathtt{p_{1}},\,\mathtt{p_{2}},\,\mathtt{q}\}$ and attributes given by $\tau(\mathtt{p_{1}})=\tau(\mathtt{p_{2}})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$ and $\tau(\mathtt{q})=\left((\mathbb{R},\,\mathtt{displacement}),\,\mathsf{s}\right)$. Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy by exchanging momentum, along with information about the current amount of momentum. Port $\mathtt{q}$ is a state port that can only exchange information about the current displacement. ∎
@@ 2.5 Bundles of port variables
A state port has one port variable called state. A power port has two additional port variables called flow and effort. Given an interface, the port variables of all its ports combined take values in a vector bundle, called the bundle of port variables associated to the interface.
@@
The state space associated to interface $I$ is given, up to isomorphism, by the Cartesian product
$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.$
To define $\mathcal{X}_{I}$ strictly, we assume that the elements of $I$ are in lexicographic order and that the Cartesian product preserves this order. ∎
@@
Given any interface $I$, each port $p\in I$ has an associated state variable $p\mathtt{.x}\in\mathcal{X}_{p}$. If $p$ is a power port, it additionally has a flow variable $p\mathtt{.f}$ and an effort variable $p\mathtt{.e}$. All port variables implicitly depend on time. As detailed in Section 6.1, the flow variable is the instantaneous rate of change of the state variable and the effort variable is the differential of a function of the state variable. Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see, e.g., [47]). The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$ yields the instantaneous power that is exchanged through port $p$.
Now we can formally define the bundle of port variables associated to an interface:
@@
Let $I$ be an interface. Based on the Boolean attribute of its ports, we can think of its associated state space $\mathcal{X}_{I}$ as the Cartesian product $\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports. Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables, and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote the flow and effort variables. The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined, up to isomorphism, as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$. Hence, $\mathcal{P}_{I}\cong\mathbb{T}\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$.
[…]
Again, to define the bundle of port variables strictly, we assume that the elements of $I$ are in lexicographic order and that the port variables are written in the same order as the ports. ∎
@@
Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has the total space (or port space)
@@
and the base space (or state space)
@@
3 Syntax
As introduced in Section 1.2, an interconnection pattern expresses how a composite system is formed from a finite number of subsystems. Syntax refers to the entirety of such expressions and their composition, i.e. their hierarchical nesting.